OpenStudy (anonymous):

A) For the uniform discrete distribution with possible values 1,2,3,...n, prove that μx = n+1/2 B) For the binomial distribution with parameters n and 0, prove that b(0;n,θ)+b(1;n,θ)+....+b(n;n,θ) = 1 C) For the binomial distribution with parameters n and θ prove that μx = nθ

OpenStudy (freckles):

what can we use for the first one

OpenStudy (freckles):

can we use \[\mu_x=x_1p_1+x_2p_2+\cdots +x_np_n\]

OpenStudy (freckles):

if so the proof should be easy for that one

OpenStudy (anonymous):

oh

OpenStudy (freckles):

\[p_i=\frac{1}{n} \text{ for all } i=1,2,3,...,n \\ \text{ and we also know } \\ 1+2+\cdots +n=\frac{n(n+1)}{2}\]

OpenStudy (freckles):

see if you can use those pieces to put your proof together

OpenStudy (freckles):

in the second one what does b(i,n, theta) mean ?

OpenStudy (anonymous):

b(0;n,θ)+b(1;n,θ)+....+b(n;n,θ) = 1

OpenStudy (anonymous):

is a #1

OpenStudy (freckles):

oh yeah sorry a is 1 b is 2 c is 3 I forgot they were lettered and not numbered

OpenStudy (freckles):

i guess b(i,n,theta) just means the probability of i happening given n the number of trials and theta being the probability of success

OpenStudy (anonymous):

ohhh

OpenStudy (freckles):

so I'm thinking \[b(i,n, \theta)= \left(\begin{matrix}n \\ i\end{matrix}\right) \theta^i(1-\theta)^{n-i}\]

OpenStudy (anonymous):

so what values do i need to plug in there?

OpenStudy (freckles):

you need to show \[\sum_{i=0}^{n} b(i, n, \theta)=1\]

OpenStudy (freckles):

that is you want to show \[\left(\begin{matrix}n \\ 0\end{matrix}\right) \theta^0(1-\theta)^{n-0}+\left(\begin{matrix}n \\ 1\end{matrix}\right) \theta^1(1-\theta)^{n-1}+\left(\begin{matrix}n \\ 2\end{matrix}\right) \theta^2(1-\theta)^{n-2} + \\ \cdots + +\left(\begin{matrix}n \\n-1\end{matrix}\right)\theta^{n-1}(1-\theta)^{n-(n-1)}+\left(\begin{matrix}n \\ n\end{matrix}\right) \theta^n(1-\theta)^{n-n}=1\]

OpenStudy (anonymous):

ohhh i got it now

OpenStudy (freckles):

and yes I know that there is a ++ that was a type-o should just be +

OpenStudy (anonymous):

no problem

OpenStudy (freckles):

like the last one should be similar to the first since they are both discrete

OpenStudy (freckles):

except i do think the third one will be a bit harder to show than the first

OpenStudy (freckles):

what i mean by similar is that you still want to use \[\mu_x=x_0p_0+x_1p_1+x_2p_2+\cdots +x_n p_n\] where \[p_i=b(i,n, \theta)=\left(\begin{matrix}n \\ i\end{matrix}\right) \theta^{i} (1-\theta)^{n-i}\]

OpenStudy (freckles):

\[\mu_x=0 \cdot p_0+1 \cdot p_1+2 \ \cdot p_2+\cdots +n \cdot p_n \\ \mu_x=1p_1+2p_2+\cdots +np_n\]

OpenStudy (anonymous):

oh i see

OpenStudy (freckles):

this might help you https://proofwiki.org/wiki/Expectation_of_Binomial_Distribution

OpenStudy (anonymous):

thanks a lot for everything

OpenStudy (freckles):

as we do see here theta+(1-theta)=1

OpenStudy (freckles):

just like there p+q=1

OpenStudy (freckles):

and no problem

OpenStudy (freckles):

i hope it helped :)

OpenStudy (anonymous):

You really did a great job!

OpenStudy (freckles):

thanks :)